# HLAT: High-quality Large Language Model Pre-trained on AWS Trainium

### v1

@@ 4.1. Training Curves
Figure 1. HLAT  training progress. (a) The training loss vs number of tokens (in billions) seen by the model during training. Gradient/Parameter norm vs number of tokens in (b)/(c), respectively. (d) The learning rate schedule vs number of tokens. The warm-up steps are 2000 iterations (about 8 billion tokens, see Section 3.2).
@@
During the training process, we monitor the training loss, as well as $l_{2}$ norm of gradients and $l_{2}$ norm of parameters for debugging training stability. Figure 1 shows the training loss over global batches, reduced over all data parallel ranks. The training loss decreases fast for the initial $\sim$250B tokens, and enters a log-linear decrease afterwards. Similar trends are observed in other LLM training (Touvron et al., 2023a, b; Geng and Liu, 2023).
@@
In Figure 1, we show the gradient $l_{2}$ norm during the training. Overall, we see that the gradient norm is stable across the training journey without divergence. Note that gradient spikes are imminent in LLM pre-training when using layer-normalization, or even RMSNorm (Takase et al., 2023), and some times due to overflow in low-precision, such as 16-bit floats. We show an assuring trend in Figure 1 even with using 16-bit floats such as BF16. Furthermore, we clip the gradient-norm to 1.0 (see Section 3.2) which also allows for magnitude stabilization. Note that sustained spikes in the gradient norm leads to training divergence due to improper weight updates, even after gradient normalization through clipping. In Figure 2, we show that the gradient spikes often last for a single step, and did not lead to training divergence. Specifically, we first track a running average ($r$) of gradient norm over a window of 20 steps to smooth out the natural fluctuations due to batching. We define occurrence of a gradient spike when the current gradient norm is higher than $r+0.1$. Next, we track the number of steps for gradient norm returning to less than $r+0.1$. Over 86%, the spike deviates from running average for only a single step.
@@
Figure 2. Number of occurrence of sustained gradient spikes vs contiguous length of appearance. Over 86%, the spike lasts for only a single step.
@@
Finally, we show the parameter $l_{2}$ norm in Figure 1. During first $\sim$250B tokens, the parameter norm increases consistently. This phase also coincides with the fast decreasing phase of training loss where model parameters converge from random initialization to a structured distribution. After that, the parameter norm consistently decrease since AdamW applies weight decay for regularization (Loshchilov and Hutter, 2019).
4.2. Hardware and System Failures
The pre-training process can be interrupted due to hardware failures, communication timeouts, etc (Borzunov et al., 2024). We monitor the node active time and frequency of training interruptions. As listed in Table 1, the overall system uptime of HLAT  pre-training is 98.81%. With automatic fault recovery mechanism in NDTL (see Section 2), the training quickly recovers from failures automatically without wasting much time. For comparison, we also listed another experimental training run (over 600 billion tokens) without automatic fault recovery, and we observed an average of 20% lower system up time.
Table 1. Percentage of system uptime with vs without automatic fault recovery.
With Fault Recovery Without Fault Recovery
98.81% 77.83%
4.3. Training Instability
@@
Initialization: We use a scaled initialization strategy for initializing model parameters. Specifically, the initial standard deviation of output layers in attention blocks and MLP layers are scaled by $1/\sqrt{2l}$ where $l$ is the layer index. Similar as discussed in (Takase et al., 2023), we found better numerical stability and convergence with smaller initial variance on deeper layers. In addition, all parameters are initialized on CPU and offloaded to Trainium.
@@
Normalization: We used tensor parallelism to shard the model parameter matrices except normalization layers. The normalization layer weights, however, are slightly different across TP ranks due to stochastic rounding. Empirically, we found the differences are small, and RMSNorm weights values are all close to 1. OLMo (Groeneveld et al., 2024) used non-parametric layer norm, which is equivalent as all weights equals 1. Both HLAT  and OLMo show similar quality as LLaMA, despite of differences in normalization weights.
Gradient Synchronization: Since Neuron SDK uses a different underlying collective library as other accelerators. It needs careful attention while writing distributed training library, such as NDTL. Mis-referencing APIs may cause training instability.
@@
Neuron Persistent Cache on Local Worker: In HLAT  training, all instances share a same file system using Amazon FSx [^8] https://aws.amazon.com/fsx/ for storing data, checkpoints, logs, etc. However, we found that storing Neuron Persistent Cache [^9] https://awsdocs-neuron.readthedocs-hosted.com/en/latest/general/arch/neuron-features/neuron-caching.html on FSx may cause communication bottleneck because those cached graphs are frequently accessed by all Trainium  devices in the cluster. Such bottleneck may lead to communication timeout and affects training stability. Therefore, we instead store Neuron Persistent Caches in file system of each local worker.
5. Evaluation
@@
Baselines: We evaluate HLAT against several open-source benchmark models. Since HLAT  structure is similar as LLaMA model, we include LLaMA-1 7B (Touvron et al., 2023a), LLaMA-2 7B (Touvron et al., 2023b), OpenLLaMA-1 7B and OpenLLaMA-2 7B (Geng and Liu, 2023). The model architecture and composition of the training data of the models being compared are listed in Table 2. OpenLLaMA-1 model is trained on RedPajama (Computer, 2023) dataset. OpenLLaMA-2 model shares same structure as OpenLLaMA-1, but is trained on a different data mixture which includes data from Falcon-RefinedWeb (Penedo et al., 2023), StarCoder (Li et al., 2023), and RedPajama (Computer, 2023).
Table 2. Comparison of model architectures and number of training tokens.
Model Name Size Sequence length Tokens
HLAT 7B 4096 1.8T
LLaMA-1 7B 2048 1T
LLaMA-2 7B 4096 2T
OpenLLaMA-1 7B 2048 1T
OpenLLaMA-2 7B 2048 1T
@@
Table 3. Evaluation of HLAT  against 4 open-source models on 7 groups of tasks describted in Section 5. For HLAT, the results are reported using the final (1.8T token) checkpoint.
Task Shots Metric OpenLLaMA-1 OpenLLaMA-2 LLaMA-1 LLaMA-2 HLAT
MMLU 5 accuracy 30.552 (3.432) 41.075 (3.611) 35.1 45.3 41.318 (3.602)
BBH 3 multiple choice grade 35.535 (1.864) 35.502 (1.861) 30.3 32.6 36.565 (1.845)
Commonsense Reasoning 0 accuracy 55.587 (1.203) 56.893 (1.195) - - 56.152 (1.194)
0 accuracy (norm) 58.411 (1.201) 61.262 (1.19) 67.3* 67.5* 59.455 (1.206)
World Knowledge 5 exact match 38.942 (0.532) 37.023 (0.52) 46.2* 48.9* 38.846 (0.534)
Reading Comprehension 0 accuracy 70.459 (0.798) 72.416 (0.782) 76.5 77.4 72.508 (0.781)
Math 8 accuracy 5.08 (0.605) 5.231 (0.613) 11.0 14.6 9.401 (0.804)
Code 0 pass@1 4.77 9.06 10.5 12.8 7.62
0 pass@10 12.83 23.58 - - 19.83
0 pass@100 23.78 40.24 36.5 45.6 34.15
@@
Evaluation Tasks: We evaluate the pre-trained model on 7 groups of tasks including both zero-shot and few-shot evaluations (Chang et al., 2023). We use Language Model Evaluation Harness (Gao et al., 2021) for natural language tasks, and use HumanEval (Chen et al., 2021) for coding tasks.
@@
  • MMLU (Massive Multitask Language Understanding) (Hendrycks et al., 2021a, b) contains 57 test tasks, spanning STEM, humanities, social sciences, and other subjects. The difficulty ranges from elementary to professional levels. The breadth of the dataset is suitable to test model’s overall problem solving and knowledge ability.
  • BIG-Bench Hard (BBH) (Suzgun et al., 2022) is a suite of 23 challenging BIG-Bench tasks (Srivastava et al., 2022), for which prior language models did not outperform the average human-rater. It can evaluate the model’s abilty on challenging tasks.
@@
  • Commonsense Reasoning consists of 6 commonly-used datasets: PIQA (Bisk et al., 2020), HellaSwag (Zellers et al., 2019), WinoGrande (Sakaguchi et al., 2021), ARC easy and challenge (Clark et al., 2018), and OpenBookQA (Geng and Liu, 2023). Those multi-choice tasks include carefully crafted riddles, puzzles, and scenarios designed to probe a model’s ability to leverage implicit knowledge, make logical inferences, and navigate the unsaid rules of our physical and social worlds.
@@
  • World Knowledge includes NaturalQuestions (Kwiatkowski et al., 2019) and TriviaQA (Joshi et al., 2017). Both tasks are designed to test model’s question-answering ability in closed book setting. The models are not provided documents that may contain information about the question, and it has to rely on information learnt or memorized in pre-training data.
@@
  • Reading Comprehension uses BoolQ (Clark et al., 2019) to test model’s open book comprehension ability. BoolQ is a question answering dataset for yes/no questions. Each example is a triplet of (question, passage, answer), with the title of the page as optional additional context. The model is required to answer the question based on the given context in passage.
@@
  • Math ability is evaluated with GSM8K (Grade School Math 8K) (Cobbe et al., 2021). GSM8K contains 8,500 grade school math problems. Both problems and answers are provided in natural language. These problems take between 2 and 8 steps to solve, which is ideal for testing basic multi-step reasoning ability of the model.
@@
  • Code evaluation uses HumanEval (Chen et al., 2021) dataset includes 164 programming problems with a function signature, docstring, body, and several unit tests. They were handwritten to ensure not to be included in the training set of code generation models.
5.1. Performance against Opensource models
We compare the performance of HLAT using the final (1.8T tokens) checkpoint with other opensource benchmarks in Table 3. For OpenLLaMA-1 and OpenLLaMA-2, we use the available pre-trained weights and evaluate using the same evaluation pipeline as our pre-trained model. For LLaMA-1 and LLaMA-2, we directly use results from corresponding papers (Touvron et al., 2023a, b). For all evaluation tasks, we adopt same hyper-parameters (number of shots) as used in (Touvron et al., 2023b). We include both average performance and variance (in the parentheses, if available). All numbers are reported in percentage.
On MMLU tasks, HLAT  performs better than OpenLLaMA-1 and LLaMA-1, and is worse then LLaMA-2. As discussed in (Touvron et al., 2023a), this may due to the difference in training datasets. We also compare the detailed performance of models on each MMLU subjects, including STEM, humanities, social science, and others (see Table 5 in Appendix). The performance is slightly worse in STEM, but similar trends is observed in OpenLLaMA-2. Comparing with LLaMA-2, HLAT  is comparable on STEM and humanities. The gaps are mainly on social science and others.
@@ 5.1. Performance against Opensource models
On BBH, Commonsense Reasoning, and World Knowledge tasks, HLAT  performs similar as OpenLLaMA-1 and OpenLLaMA-2 models. By diving deep into performance on each individual task (see Table 5 in Appendix), HLAT  excels in 19/29 tasks as compared with OpenLLaMA-1, and 15/29 tasks compared with OpenLLaMA-2. Both HLAT  and OpenLLaMA models have some gaps with LLaMA-1 and LLaMA-2 models. This might be due to using different evaluation libraries and slightly difference in datasets. Specifically, the gaps in Commonsense Reasoning and World Knowledge are mainly from OpenBookQA and TriviaQA, respectively. On TriviaQA, The results of LLaMA models (Touvron et al., 2023b) are reported on wikipedia subset; whereas HLAT  and OpenLLaMA are evaluated on entire dataset.
@@
On Math problems (GSM8K), HLAT  performs significantly better than OpenLLaMA-1 and OpenLLaMA-2. As will be discussed in the next section, HLAT  has a big improvement of Math ability in later training stage.
@@
On Coding problems, HLAT  performs better than OpenLLaMA-1, comparable with LLaMA-1, and worse than OpenLLaMA-2 and LLaMA-2. First, for OpenLLaMA-1, the tokenizer merges consecutive spaces which negatively affects the coding performance, as it eliminates important information such as indentation and line breaks. This issue is subsequently fixed in OpenLLaMA-2, which explains its better performance. Besides, OpenLLaMA-2 is trained with additional code data from StarCoder which also contributes to performance improvement.
5.2. Intermediate Model Performance
Figure 3. Intermediate model performance with number of seen tokens. (a) Commonsense Reasoning, (b) Math, (c) World Knowledge. We observe different trends of learning curves for different tasks. Detailed results of other tasks are in Figure 5.
@@ 5.2. Intermediate Model Performance
During the model training, we also evaluate the intermediate checkpoints every 200 billion tokens. On most benchmarks, the performance improves steadily, and correlates with the training loss of the model (see Figure 1). Detailed evaluation results are shown in Appendix B (Table 6 and Figure 5).
We found that for different tasks, the model converges at different rates. Figure 3 plots the model performance on three sets of evaluation tasks with respective to number of seen training tokens (in billions).
@@
For Commonsense Reasoning, the model accuracy improves quickly at beginning of training, and starts to saturate at later training stages. This is similar as the trends observed in other LLM model trainings (Touvron et al., 2023a; Groeneveld et al., 2024).
@@
For World Knowledge task shown in Figure 3, the performance increases almost linearly with number of training tokens. Since this is a closed book test and mainly evaluates the model’s ability of memorizing facts in pre-training data, the model seems to consistently improve its ability on this domain with more training steps and epochs. In addition, we also tested if the trending is related to number of shots used in evaluation. The trends are very similar for zero-shot, 3-shot, and 5-shot tests.
@@
Those observations indicate the necessity of a set of evaluation tasks covering a wide range of domains for LLM pre-training. A single validation set or evaluation tasks from narrow domains may not comprehensively reflect the actual over- or under-fitting of the model for general downstream tasks.
5.3. Truthfulness and Bias
Table 4. Model Truthfulness and Bias evaluation.
Dataset Domains Metric OpenLLaMA-1 OpenLLaMA-2 LLaMA-1 HLAT
CrowS-Pairs age pct_stereotype 62.93 (4.997) 57.399 (5.1) 70.1 61.819 (5.002)
physical appearance pct_stereotype 66.667 (5.495) 73.611 (5.177) 77.8 62.5 (5.62)
race/color pct_stereotype 47.022 (2.226) 50.282 (2.256) 57.0 51.907 (2.271)
disability pct_stereotype 69.499 (5.702) 66.48 (5.757) 66.7 65.676 (5.893)
gender pct_stereotype 54.925 (2.756) 57.887 (2.742) 70.6 61.788 (2.696)
nationality pct_stereotype 48.969 (3.174) 45.502 (3.174) 64.2 51.273 (3.22)
sexual orientation pct_stereotype 77.207 (4.38) 80.946 (4.1) 81.0 82.607 (3.974)
religion pct_stereotype 72.652 (4.179) 70.043 (4.257) 79.0 68.723 (4.305)
socioeconomic pct_stereotype 68.649 (3.348) 72.309 (3.225) 71.5 66.88 (3.391)
TruthfulQA - multiple choice 1 23.133 (1.476) 22.644 (1.465) - 23.623 (1.487)
multiple choice 2 35.141 (1.355) 34.576 (1.348) - 37.188 (1.349)

### v2

@@ IV-A Training Curves
Figure 1: HLAT training progress. (a) The training loss vs number of tokens (in billions) seen by the model during training. Gradient/Parameter norm vs number of tokens in (b)/(c), respectively. (d) The learning rate schedule vs number of tokens. The warm-up steps are 2000 iterations (about 8 billion tokens, see Section III-A).
@@
During the training process, we monitor the cross entropy training loss, as well as $l_{2}$ norm of gradients and $l_{2}$ norm of parameters for debugging training stability. Figure 1 shows the training loss over global batches, reduced over all data parallel ranks. The training loss decreases rapidly over the initial $\sim$250 billion tokens, and enters a log-linear decrease afterwards. Similar trends are observed in other LLM pre-training [8, 9, 29].
In Figure 1, we show the gradient $l_{2}$ norm during the training. Overall, we see that the gradient norm is stable across the training journey without divergence. Note that gradient spikes are common in LLM pre-training when using layer-normalization, or even RMSNorm [30], and sometimes due to overflow in low-precision, such as 16-bit floats. We show an assuring trend in Figure 1 even with using 16-bit floats.
@@
Note that sustained spikes in the gradient norm leads to training divergence due to improper weight updates, even after gradient normalization through clipping (see Section III-A). In Figure 2, we show that the gradient spikes often last for a single step, and did not lead to training divergence. Specifically, we first track a running average ($r$) of gradient norm over a window of 20 steps to smooth out the natural fluctuations due to batching. We define occurrence of a gradient spike when the current gradient norm is higher than $r+0.1$. We then track the number of steps for gradient norm returning to less than $r+0.1$. Over 86%, the spike deviates from running average for only a single step.
@@
Figure 2: Number of occurrence of sustained gradient spikes vs contiguous length of appearance. Over 86%, the spike lasts for only a single step.
@@
Finally, we show the parameter $l_{2}$ norm in Figure 1. During first $\sim$250B tokens, the parameter norm increases consistently. This phase also coincides with the fast decreasing phase of training loss where model parameters converge from random initialization to a structured distribution. After that, the parameter norm consistently decrease since AdamW applies weight decay for regularization [31].
IV-B Hardware and System Failures
Pre-training on a cluster with thousands of accelerators often faces hardware-level errors that interrupts the training process. Those errors could be due to malfunction of AWS Trainium chips or host machines, network communication timeouts, and so on [32]. Manually restarting the pre-training from last-saved checkpoints demands significant developer time and causes low system uptime. For example, we performed an experimental training run (over 600 billion tokens) without automatic fault recovery, and we observed an average system uptime of 77.83%. In HLAT pre-training, we enabled automatic fault recovery mechanism in NxDT (see Section II), which automatically replaces the faulty nodes and restart training from latest checkpoints. The overall system uptime of HLAT pre-training is then improved 20% to 98.81%.
IV-C Training Convergence and Instability
@@
Initialization: We use a scaled initialization strategy for initializing model parameters. Specifically, the initial standard deviation of output layers in attention blocks and MLP layers are scaled by $1/\sqrt{2l}$ where $l$ is the layer index. Similar as discussed in [30], we found better numerical stability and convergence with smaller initial variance on deeper layers.
@@
Normalization: We used tensor parallelism to shard the model parameter matrices except normalization layers. For HLAT-7B, the normalization layer weights, however, are slightly different across TP ranks due to stochastic rounding. Empirically, we found the differences are small, and RMSNorm weights values are all close to 1. Note that with standard mixed precision strategy, we do not see such weight difference for HLAT-70B.
Checkpoint Averaging: For HLAT-70B, the final checkpoint used for evaluation is an average of last two checkpoints in the training process. Similar as [10, 33], we found averaging the last two checkpoints provides better performance than each single checkpoint. We provide more detailed discussion in Section V-D.
@@
Neuron Persistent Cache on Local Worker: In HLAT pre-training, all instances share a same file system using Amazon FSx [^8] https://aws.amazon.com/fsx/ for storing data, checkpoints, logs, etc. However, we found that storing Neuron Persistent Cache [^9] https://awsdocs-neuron.readthedocs-hosted.com/en/latest/general/arch/neuron-features/neuron-caching.html on FSx may cause communication bottleneck because those cached graphs are frequently accessed by all Trainium devices in the cluster. Such bottleneck may lead to communication timeout and affects training stability. Therefore, we instead store caches in file system of each local worker.
V Evaluation
@@
Baselines: We evaluate HLAT against several open-source benchmark models. Since HLAT structure is similar as LLaMA model, we include LLaMA-1 (7B, 13B, 65B) [8], LLaMA-2 (7B, 70B) [9], OpenLLaMA-1 (7B, 13B) and OpenLLaMA-2 (7B) [29]. The model architecture and composition of the training data of the models being compared are listed in Table I. OpenLLaMA-1 model is trained on RedPajama [21] dataset. OpenLLaMA-2 model shares same structure as OpenLLaMA-1, but is trained on a different data mixture which includes data from Falcon-RefinedWeb [34], StarCoder [35], and RedPajama [21].
TABLE I: Model architectures comparison.
Model Sizes Sequence length
OpenLLaMA1 7B, 13B 2048
OpenLLaMA2 7B 2048
LLaMA1 7B, 13B, 33B, 65B 2048
LLaMA2 7B, 13B, 70B 4096
HLAT 7B, 70B 4096
@@
TABLE II: Evaluation of HLAT against 4 open-source models on 6 groups of tasks described in Section V. Numbers in the parentheses represent standard deviation, if available.
Model Size MMLU CR WK RC Math Code Average
- - accuracy accuracy exact match accuracy accuracy pass@1 pass@10 -
OpenLLaMA-1 7B 30.5 58.4 40.6 70.5 5.2 4.5 13.4 41.2
OpenLLaMA-2 7B 41.1 61.3 37.9 72.4 6.8 9.7 25 44.9
LLaMA-1 7B 35.1 63.5 43.6 76.5 11 10.5 21.3 47.4
LLaMA-2 7B 45.3 64 45.2 77.4 14.6 12.2 25.2 49.2
HLAT-7B 7B 41.3 (3.6) 59.5 (1.2) 38.8 (0.5) 72.5 (0.8) 9.4 (0.8) 7.6 19.8 44.6
OpenLLaMA-1 13B 43.5 62 45.9 72.3 8.3 7 17 47.1
LLaMA-1 13B 46.9 65.3 49.7 78.1 17.8 15.8 22.6 53.1
LLaMA-2 13B 45.3 66.3 50.5 81.7 28.7 18.3 30.5 54.6
LLaMA-1 33B 57.8 68.9 54.6 83.1 35.6 21.7 38.4 59.2
LLaMA-1 65B 63.4 69.8 57 85.3 50.9 23.7 - 62.1
LLaMA-2 70B 68.9 70.7 59 85 56.8 30.5 59.4 64.7
HLAT-70B 70B 65.1 (3.4) 67.3 (1.2) 54.5 (0.6) 82.6 (0.7) 48.5 (1.4) 21.4 57.9 60.8
@@
Evaluation Tasks: We evaluate HLAT against baselines on 7 groups of tasks including both zero-shot and few-shot tasks [36]. We use HumanEval [37] for coding tasks, and Language Model Evaluation Harness [38] for others.
@@
Massive Multitask Language Understanding (MMLU) [15, 39] contains 57 tasks, spanning STEM, social sciences, humanities, and other subjects. The difficulty ranges from elementary to professional levels. The breadth of the dataset tests model’s overall problem solving and knowledge ability.
@@
Commonsense Reasoning (CR) consists of 6 datasets: PIQA [40], HellaSwag [41], WinoGrande [42], ARC easy and challenge [43], and OpenBookQA [29]. These multi-choice tasks include carefully crafted riddles, puzzles, and scenarios designed to probe a model’s ability to leverage implicit knowledge, make logical inferences, and navigate the rules of physical and social worlds.
@@
World Knowledge (WK) includes NaturalQuestions [44] and TriviaQA [45]. Both tasks are designed to test model’s question-answering ability in closed book setting. The models are not provided documents that may contain information about the question, and it has to rely on information learnt or memorized in pre-training data.
@@
Reading Comprehension (RC) uses BoolQ [46] to test model’s open book comprehension ability. BoolQ is a question answering dataset for yes/no questions. Each example is a triplet of (question, passage, answer), with the title of the page as optional additional context. The model is required to answer the question based on the given context in passage.
@@
Math ability is evaluated with GSM8K (Grade School Math 8K) [47]. GSM8K contains 8,500 grade school math problems. Both problems and answers are provided in natural language. These problems take between 2 and 8 steps to solve, which is ideal for testing basic multi-step reasoning ability.
@@
Code evaluation uses HumanEval [37] dataset including 164 programming problems with a function signature, docstring, body, and several unit tests. They were handwritten to ensure not present in the training set of the models.
V-A Performance against open-source Models
We compare the performance of HLAT with other open-source benchmarks in Table II. The numbers are reported in percentage and for HLAT results, we include both mean and standard deviation (in the parentheses, if available). We also report an average score over all tasks in the last column.
HLAT-7B performs better than OpenLLaMA-1 and is on-par with OpenLLaMA-2. Both HLAT-7B and OpenLLaMA models have some gap with LLaMA-1 and LLaMA-2, which is likely due to the training data quality. Even though the data composition of RedPajama-1T is similar as those used in LLaMA-1, the data cleaning pipeline and final data quality are different, which therefore affects the model performance [48]. For HLAT-70B, we use the same training dataset as the 7B model for consistency. Although there is no OpenLLaMA baseline for a fair comparison, HLAT-70B performs better than LLaMA-1 and LLaMA-2 models of smaller sizes. The model performance gap with LLaMA-1 (65B) and LLaMA-2 (70B) is also smaller than those on 7B models. We acknowledge the lack of effort on data quality improvement, but our main goal is to showcase the effectiveness and efficiency of AWS Trainium.
On MMLU (5-shot), both HLAT models perform better than OpenLLaMA-1 and LLaMA-1 models of similar size. The performance is slightly worse than LLaMA-2 family of models, likely due to the difference in training dataset size and composition [8].
@@ V-A Performance against open-source Models
On Commonsense Reasoning (0-shot) and World Knowledge (5-shot), HLAT-7B performs similar to OpenLLaMA-1 and OpenLLaMA-2 models. By diving deep into performance on each individual task, HLAT-7B excels in 19/29 tasks as compared with OpenLLaMA-1, and 15/29 tasks compared with OpenLLaMA-2. Both HLAT and OpenLLaMA models have some gaps with LLaMA-1 and LLaMA-2 models, which may be due to the training set quality. Nevertheless, the gap ($\sim 3\%$) is consistent on 7B and 70B models.
@@
On Math problems (GSM8K, 8-shot), HLAT-7B performs significantly better than OpenLLaMA-1 and OpenLLaMA-2. As will be discussed in the next section, HLAT-70B has a big improvement of Math ability in later training stage. HLAT-70B performs similar as LLaMA1-65B, and we observed significant improvement in upsampling training stage.
@@
On Coding problems, both HLAT-7B and HLAT-70B perform comparable with LLaMA-1. HLAT-7B performs better than OpenLLaMA-1 and worse than OpenLLaMA-2 and LLaMA-2. First, for OpenLLaMA-1, the tokenizer merges consecutive spaces which negatively affects the coding performance, as it eliminates important information such as indentation and line breaks. This issue is subsequently fixed in OpenLLaMA-2, which explains its better performance. Besides, OpenLLaMA-2 is trained with additional code data from StarCoder which also contributes to performance improvement.
V-B Intermediate Model Performance
Figure 3: Intermediate model performance with number of seen tokens for HLAT-7B.
Figure 4: Intermediate model performance with number of seen tokens for HLAT-70B.
@@ V-B Intermediate Model Performance
During the model training, we also evaluate the intermediate checkpoints about every 200 billion tokens. Figure 3 and Figure 4 show the model performance of HLAT-7B and HLAT-70B with respect to number of seen training tokens (in billions), respectively. On most benchmarks, the performance improves steadily, and correlates with the training loss.
@@
We found that for different tasks, the model converges at different rates. For Commonsense Reasoning, the model accuracy improves quickly at beginning of training, and starts to saturate at later training stages. This is similar as the trends observed in other LLM model trainings [8, 49].
@@
For World Knowledge task shown in Figure 3, the performance increases almost linearly with number of training tokens. Since this is a closed book test and mainly evaluates the model’s ability of memorizing facts in pre-training data, the model seems to consistently improve its ability on this domain with more training steps and epochs. In addition, we tested if the trending is related to number of shots used in evaluation. It turns out that the trends are very similar for zero-shot, 3-shot, and 5-shot tests.
@@
Those observations indicate the necessity of a set of evaluation tasks covering a wide range of domains for LLM pre-training. A single validation set or evaluation tasks from narrow domains may not fully reflect the actual over- or under-fitting of the model for general downstream performance.
V-C Upsampling
TABLE III: Upsampling dataset composition for HLAT-70B.
Datasets Size Percentage
(billions of tokens)
Web Data Wikipedia[21] 90 35.47%
C4[21]
Domain Specific StackExchange 104.7 41.27%
Arxiv[21]
Open-Web-Math[23]
PeS2o[22]
Code Github[21] 59 23.26%
Total - 253.7 15.16%
During HLAT-70B training, we upsampled the training dataset in last 400B tokens. Specifically, we use 35.47% web data, 41.27% math data, and 23.26% coding data with more details listed in Table III. In Figure 4, we plot the evaluation performance of HLAT-70B with seen training tokens. In upsampling training stage, that is, after 1400B tokens, we observe significant model performance improvement over math, coding, and MMLU performance. It improved math by 10% and coding by 5%. This is consistent with the findings in LLaMA-3 [10], where the researchers found significant improvement of LLaMA-3 8B model on math problems. However, they mentioned such method did not help much for 405B models. Our experiment fills the model size gap, and shows that upsampling still helps for a 70B model.
V-D Checkpoint Averaging
For HLAT-70B, we average the last two checkpoints used in pre-training to generate a checkpoint for final evaluation [10, 33]. Table IV compares the individual checkpoints with 1740B training tokens and 1800B training tokens, as well as the averaged checkpoints. The averaged checkpoints outperforms individual checkpoints on average performance, as well as some individual tasks.
TABLE IV: Evaluation of HLAT-70B with individual and averaged checkpoints.
Checkpoint MMLU RC WK CR Math Code Avg.
1740B 64.5 67.5 54 83.1 47.3 18.3 60.3
1800B 64.2 66.9 54 82.1 47.2 21.8 60.2
Average 65.1 67.3 54.5 82.6 48.5 21.4 60.8
V-E Truthfulness and Bias